OpenStudy (anonymous):

The measure of an an angle is 4/5 the measure of its complement. Find the measure of each angle.. Help

OpenStudy (barrycarter):

Ummm, that's not quite what the problem says, @campbell_st -- the measure if 4/5th of the compliment, not 4/5th of 90 degrees

OpenStudy (anonymous):

yap

OpenStudy (campbell_st):

ok then its x + 4/5 x = 90....

OpenStudy (campbell_st):



OpenStudy (anonymous):

thanks Everyonee!..

OpenStudy (anonymous):

that's it\

OpenStudy (anonymous):

Huh?

OpenStudy (campbell_st):

its simply 9x/5 = 90 solve for x.... which will be the larger angle

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Wait um Im confused in what to do to find x

OpenStudy (campbell_st):

multiply both sides by 5 9x = 450 divide both sides by 9

OpenStudy (anonymous):

Wghere did you get 9x from? Can you do it step by step?

OpenStudy (campbell_st):

if you have 1 + 4/5 = 1 4/5 or 9/5 as an improper fractor

OpenStudy (anonymous):

Thanks again I understand